# An Empirical Analysis on the Use and Reporting of National Security Letters

### v3

@@ A.1 List of Transparency Report Companies
A.2 Number of NSLs per Company
@@
Table 2: Statistics on NSLs published by companies
Appendix B Nondisclosure duration time
B.1 Bias characterization and numerical experiments
@@
Figure 6: This graph shows schematically how we interpret the NSL letters as observations drawn from some distribution.
@@
$P_{observed}(a,b)\propto\mathds{1}\{a_{1}+b_{1}\leq k\}\cdot P_{published}({published}|a=a_{1},b=b_{1})\cdot P_{lifted}(b=b_{1}|a=a_{1})$
@@
Probability of a letter to be published is constant
$P_{published}({published}|a=a_{1},b=b_{1})=p_{pub}$
@@
Gag order duration time is uniformly distributed
$P_{lifted}(b=b_{1}|a=a_{1})\sim U[0,N]$
@@
B.2 True mean estimation
@@
$P_{observed}(a=a_{1})=\sum_{b\in[0,N]}P_{observed}(a=a_{1},b=b_{1})=\frac{k-a_{1}}{N}\cdot p_{pub}$
@@
$P_{observed}(a=a_{1})=Bernoulli(\frac{k-a_{1}}{N}\cdot p_{pub})$
@@
$E[B(n_{a_{1}},\frac{k-a_{1}}{N}\cdot p_{pub})]=\frac{k-a_{1}}{N}\cdot p_{pub}\cdot n_{a_{1}}$
@@
$N_{a_{1}}=\frac{k-a_{1}}{\#obs_{a_{1}}}\cdot p_{pub}\cdot n_{a_{1}}$

### v4

@@ B.1. List of Transparency Report Companies
B.2. Number of NSLs per Company
@@
Appendix C Nondisclosure duration time
C.1. Bias characterization and numerical experiments
@@
Figure 6. This graph shows schematically how we interpret the NSL letters as observations drawn from some distribution.
[…]
[This graph shows schematically how we interpret the NSL letters as observations drawn from some distribution.]()
@@
(1) $\begin{split}P_{observed}(a,b)&\propto\mathds{1}\{a_{1}+b_{1}\leq k\}\\ &\cdot P_{published}({published}|a=a_{1},b=b_{1})\\ &\cdot P_{lifted}(b=b_{1}|a=a_{1}).\end{split}$
@@
Probability of a letter to be published is constant:
$P_{published}({published}|a=a_{1},b=b_{1})=p_{pub}.$
@@
Gag order duration time is uniformly distributed:
$P_{lifted}(b=b_{1}|a=a_{1})\sim U[0,N].$
@@
C.2. True mean estimation
@@
(2) $\begin{split}P_{observed}(a=a_{1})&=\sum_{b\in[0,N]}P_{observed}(a=a_{1},b=b_{1})\\ &=\frac{k-a_{1}}{N}\cdot p_{pub}.\end{split}$
@@
$P_{observed}(a=a_{1})=Bernoulli(\frac{k-a_{1}}{N}\cdot p_{pub}).$
@@
$E[B(n_{a_{1}},\frac{k-a_{1}}{N}\cdot p_{pub})]=\frac{k-a_{1}}{N}\cdot p_{pub}\cdot n_{a_{1}}.$
@@
$N_{a_{1}}=\frac{k-a_{1}}{\#obs_{a_{1}}}\cdot p_{pub}\cdot n_{a_{1}}.$